A chemistry student needs 600 ml of a 6% solution. How much 10% and 4% solutions need to be mixed?
Answer by wilft1(217)

this one is solved with 2 equations
x + y = 600
10x + 4y = 6
multiply 600 by 6
x + y = 600
10x + 4y = 3600
multiply the top line by -4
-4x - 4y = -2400
10x + 4y = 3600
add the two lines together
6x = 1200
divide both sides by 6
x = 200
now that we know the value for x, plug that into the original equation
10(200) + 4y = 3600
2000 + 4y = 3600
4y = 1600
y = 400
x = 200, y = 400
so you will need 200 ml of the 10 percent solution and 400 ml of the 4 percent solution